### v1

@@ VII Conclusion
We investigate the enhancement of the resonant absorption of the electromagnetic field by two-level absorbers (atoms or molecules) near plasmonic nanoparticles, taking into account the spontaneous and collective emission from absorbers to the localized plasmon resonance (LPR) mode of the particle. We derive quantum equations for the particle and the resonant absorber polarisation and the absorbing state populations. We analyze the structure of equations and describe the physical mechanisms of the stationary absorption and the particle-absorber interaction.
@@
We find that the absorption enhancement by the LPR is possible under certain conditions. The absorption suppression is possible as well due to the spontaneous, collective emissions of absorbers to the LPR mode and the energy transfer from absorbers to the particle. We find conditions necessary for the absorption enhancement. One condition is a fast, of the pico-second time scale, de-population of the excited absorber states, for example, by some non-radiative processes.
@@
We find conditions and investigate the bistability in the resonant absorption near the metal nanoparticle. The bistability is caused by the collective emission and the nonlinear interaction between the particle and absorbers through the LPR near field. The bistability appears at a large number of absorbers near the nanoparticle and possible with the absorption enhancement or suppression.
@@
Results can be used everywhere, when the resonant absorption near metal nanoparticles is used, for example, for the plasmon-stimulated photo-catalysis Zhang et al. (2013) or in the plasmonic photo-voltaic Jang et al. (2016).
Acknowledgements.
@@
Polarizability $\alpha_{p}$ of the spherical nanoparticle of the radius $r_{p}$ in the electric field with the wavelength $\lambda$, in the dielectric environment with the refractive index $\eta$ is Protsenko and Uskov (2012)
@@
where $\varepsilon_{p}$ is the dielectric function of the material of the nanoparticle and the factor
@@
The nanoparticle material dielectric function $\varepsilon_{p}(\lambda,r_{p})$ takes into account the losses due to collisions of the nanoparticle electrons with the nanoparticle boundary Protsenko and Uskov (2012); Uskov et al. (2014)
@@
[…]

### v2

@@ 7 Conclusion
We study the enhancement of the resonant absorption of the electromagnetic field by two-level absorbers (atoms or molecules) in the vicinity of plasmonic nanoparticles, taking into account the spontaneous and collective emission from the absorbers to the localised plasmon resonance (LPR) mode of the particle. We derive quantum equations for the polarisation of the particle and the resonant absorber, and for the populations of absorbing states. We analyse the structure of the equations and describe the physical mechanisms of the stationary absorption and the particle-absorber interaction.
@@
We find that absorption enhancement by LPR is possible under certain conditions. Absorption suppression is also possible due to individual spontaneous and collective emissions from absorbers to the LPR mode and energy transfer from absorbers to the particle. We find conditions necessary for absorption enhancement. One condition is a fast, on the picosecond time scale, depopulation of the excited absorber states, e.g. by some non-radiative processes.
@@
We find the conditions and study the bistability in the resonant absorption near the metal nanoparticle. The bistability is caused by the collective emission and the nonlinear interaction between the particle and the absorbers through the LPR near field. The bistability occurs with a strong external field excitation, a large number of absorbers in the vicinity of the nanoparticle, and is possible with absorption enhancement or suppression.
@@
The results can be applied wherever resonant absorption near metal nanoparticles is used, for example in plasmon-stimulated photocatalysis [8] or plasmonic photovoltaics [7].
Acknowledgments
@@
The polarizability $\alpha_{p}$ of a spherical nanoparticle of the radius $r_{p}$ in an electric field of wavelength $\lambda$, in a dielectric medium of refractive index $\eta$ is [35]
@@
where $\varepsilon_{p}$ is the dielectric function of the nanoparticle material and the factor
@@
The nanoparticle material dielectric function $\varepsilon_{p}(\lambda,r_{p})$ taking into account the losses due to collisions of the nanoparticle electrons with the nanoparticle boundary [35, 36] is
@@
The expression (36) with $r_{c}=0$ is a Drude dielectric function. For silver $\varepsilon_{0}=5$, $\lambda_{p}=0.14$ mkm, $\lambda_{c}=32$ mkm [67]. The term $\sim r_{c}=A_{0}v_{F}\lambda_{c}/2\pi c$ describes the correction to the dielectric function due to collisions of electrons with the nanoparticle surface. The factor $A_{0}=0.7$, $v_{F}\approx 1.3\cdot 10^{6}$ m/s is the Fermi velocity of the electron, $c$ is the speed of light in vacuum [35]. Radiation losses increase with the nanoparticle radius $r_{p}$ while the losses due to collisions with the nanoparticle surface decrease with $r_{p}$. This, there is an optimum $r_{p}$, when the nanoparticle LPR has the maximum quality factor $Q_{LPR}$. Estimates show that the maximum $Q_{LPR}\approx 29$, when $r_{p}=r_{p}^{opt}=12$ nm. The LPR wavelength $\lambda_{LPR}\approx 408$ nm and the LPR half-width $\Gamma\approx 80$ THz. The absolute value of the maximum resonant polarisability of the nanoparticles is $\alpha_{p}=|\alpha_{p}(\lambda_{LPR},r_{p}^{opt})|\approx 3\cdot 10^{-5}$ mkm3.